NvidiaIntely (nvidiaintely):

How are the coordinates of the new point found if it is rotated 90° counterclockwise? How many degrees is that equivalent to if the rotation is clockwise?

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

360-90 = 270

NvidiaIntely (nvidiaintely):

How are the coordinates of the new point found if it is rotated 180° counterclockwise? How many degrees is that equivalent to if the rotation is clockwise?

OpenStudy (will.h):

am too bored seriously and too lazy i will just watch Kevin solving problems

NvidiaIntely (nvidiaintely):

lol

OpenStudy (kevin):

okay... for the first question we should find the matrix rotation \[M = \frac{ \cos180 \times -\sin180 }{ \sin 180\times \cos 180 }\] \[= \left[\begin{matrix}-1 & 0 \\ 0 & -1\end{matrix}\right]\] so, for any (x,y) rotated by 180 degree would be \[(x,y) \times \left[\begin{matrix}-1 & 0 \\ 0 & -1\end{matrix}\right]\] = (-x, -y)

OpenStudy (will.h):

I wonder what's he typing. i never seen him typing like that before

OpenStudy (kevin):

I was typing the equation will.... lol

OpenStudy (will.h):

lol i will just watch.. looks pretty nice work keep going mate

OpenStudy (kevin):

wait,, it's 90 degree

OpenStudy (kevin):

OMG I'm writing again o_O

OpenStudy (will.h):

i don't think you needed to use all that tri functions

OpenStudy (kevin):

You have more simple method?

OpenStudy (will.h):

well there's simple form 90 degrees counterclockwise (x,y) --> (-y,x) that's a constant fact you don't need to prove anything

OpenStudy (kevin):

lol.. Honestly, my teacher never teach me like that xD He just give me the matrix rotation formula.. Thx anyway will!

OpenStudy (will.h):

lol sure np your teacher loves to show off lol

OpenStudy (kevin):

xD